OpenStudy (anonymous):

how do i calculate this no answers just how to get it i give medals plzz help me asap!!!!!!!

OpenStudy (anonymous):

Pies are ordered from the Pizza Palace according to their diameter. Large:sixteen inches Medium:fourteen inches Small:twelve inches How many more square inches are in the large pizza than in the small pizza? twelve point five six square inches forty-seven point one zero square inches eighty-seven point nine two square inches three hundred fifty-one point six eight square inches

OpenStudy (anonymous):

Remember the formula for the are of circle: \[A = pR^2\] SO the area of large one: \[A = \pi * r^2 = \pi * (16)^2 \approx 804.25\] and the area of small area: \[A = \pi \times (12)^2 \approx 452.39\] A_largepie - A-small pie = 804.25 - 452.39 = 351.68in^2

OpenStudy (anonymous):

area but yeah lol

OpenStudy (anonymous):

thanks everyone

OpenStudy (anonymous):

\(\LARGE@adjax\)